OpenStudy (anonymous):

At Water World the admission fee is $17.00, and you can rent an inner tube for $0.75 per hour. You can use the formula P = 17.00 + 0.75 h to find the total cost for admission plus renting an inner tube. a. Solve the formula P = 17.00 + 0.75 h for h. Explain your steps for solving the formula for h. b. For how many hours can you rent an inner tube if you only have $20.75 to spend? Explain your method for finding the answer

OpenStudy (anonymous):

P = 17.00 + 0.75 H - 17 from each side P - 17 = 0.75H divide each side by 0.75 H = (p-17)/0.75 H = (20.75 - 17)/0.75 H = 3.75/0.75 H = 5 (This was on wikianswers, but I think you already knew that.)